# Causal Classification of Pathological Misner-Type Spacetimes

### v2

@@ 5.1.2. Relation in the $4$-dimensional case
Ultimately, these results lead us to the following
Conjecture 5.5.
The $4$-dimensional pseudo-Schwarzschild spacetime and Misner space are isocausal.
@@ 6. Pseudo-Reissner-Nordström spacetime
By performing a Wick rotation $\theta\rightarrow i\theta$ on the Reissner-Nordström metric, changing the range of $\theta$ from $[0,\pi]$ to $[0,+\infty]$, flipping the signature, and identifying $t+P$ with $t$, we obtain the manifold
@@
Similar to the pseudo-Schwarzschild manifold, the $1$-sphere $S^{1}$ (with $0\leq t<t+\alpha$) is topologically equivalent to an interval with its endpoints identified. The surface $H_{r}^{2}$ is $2$-dimensional and corresponds to the upper sheet of a two-sheeted spacelike hyperboloid. The coordinate $\theta$ can take any positive value, while $\phi$ is periodic and ranges from $0\leqslant\phi<2\pi$. The real constants $m$ and $q$ represent mass and electric charge, respectively, and are assumed to be positive. This new metric can also be derived by performing a Wick rotation and a signature change on the Schwarzschild metric, followed by the substitution
@@
In the case of the Reissner-Nordström spacetime, matter is present—at least in the form of an electrostatic field—so it is not a vacuum solution of Einstein’s field equations. A vacuum solution requires both the Einstein tensor and the stress-energy tensor to vanish identically, meaning that no matter or non-gravitational fields are present. In contrast, the Reissner-Nordström solution includes a non-zero electromagnetic field, which contributes to the stress-energy tensor. As such, it is classified as an electrovacuum solution of the Einstein–Maxwell equations, where the electromagnetic field is sourced by a vector potential. In mathematical terms, when the Reissner-Nordström metric is written with signature ($+$)(-$+$)( - + + + ), it satisfies Einstein’s equations in the form $R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=\kappa T_{\mu\nu}$, where $T_{\mu\nu}$ is the energy–momentum tensor associated with the corresponding electromagnetic field.
Our construction of the pseudo-Reissner-Nordström spacetime involves performing a Wick rotation on the angular coordinate, $\theta\rightarrow i\theta$, resulting in a hyperbolic geometry in the angular sector. This transformation leaves the time coordinate unchanged, and the electromagnetic field strength $F=dA$ formally retains its original expression. However, in our construction, the Wick rotation is followed by a sign reversal of the metric. This change flips the sign on the right-hand side of the Einstein field equations while leaving the left-hand side invariant. Specifically, if we first apply the transformation $\theta\to i\theta$, and then reverse the overall sign of the metric, we once again obtain a Lorentzian metric with signature ($+$)(-$+$)( - + + + ), but differing from the original by an overall sign. Under this transformation, the field strength $F_{\mu\nu}$ remains unchanged, while the mixed-index version $F_{\mu}{}^{\nu}$ acquires a minus sign. Similarly, the contraction $g_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}$ changes sign. As a result, the right-hand side of the Einstein field equations changes sign, while the left-hand side remains the same.
Consequently, the resulting spacetime is no longer a solution of the Einstein–Maxwell equations. The energy–momentum tensor corresponding to the pseudo-Reissner-Nordström metric violates the weak energy condition: the gravitational source now corresponds to a form of ‘exotic matter’ with negative energy density. This situation differs significantly from the case of the pseudo-Schwarzschild metric, where the original stress-energy tensor was zero—and remains zero after the transformation. The pseudo-Schwarzschild metric, therefore, remains a valid vacuum solution to Einstein’s equations, whereas the pseudo-Reissner-Nordström metric is no longer a solution of the Einstein–Maxwell equations. This distinction is critical because it differentiates the pseudo-Reissner-Nordström spacetime from the Misner and pseudo-Schwarzschild spacetimes, both of which are vacuum solutions. What makes the pseudo-Reissner-Nordström spacetime particularly surprising and significant as a member of the causality-violating spacetime family is its non-vacuum character.
@@
We regard $\theta$ and $\phi$ being constant and first compute the tortoise coordinate $r*$ by setting
@@
[…]
$=(2m^{2}-q^{2})\cdot\tan^{-1}(\frac{r-m}{\sqrt{q^{2}-m^{2}}})-m\cdot\log(-2mr-q^{2}+r^{2})-r+c,$
@@
$ds^{2}$ $=-(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)\,(d(\nu-r*))^{2}+(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sinh^{2}(\theta)d\phi^{2})$
[…]
$+\,(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sinh^{2}(\theta)d\phi^{2})$
$=-(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)\,[(\frac{1}{(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{2}})dr^{2}-2\cdot 1\cdot\frac{1}{(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)}drd\nu+1d\nu^{2}]$
$+\,(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sinh^{2}(\theta)d\phi^{2})$
$=-(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{-1}dr^{2}+2drd\nu-(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)d\nu^{2}$
$+\,(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sinh^{2}(\theta)d\phi^{2})$
$=-(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)d\nu^{2}+2drd\nu+r^{2}(d\theta^{2}+\sinh^{2}(\theta)d\phi^{2})$.
@@
The Riemann curvature tensor for the $2$-dimensional pseudo-Reissner-Nordström spacetime can be obtained using the same approach as in Section 4 for the pseudo-Schwarzschild case. The scalar curvature then follows immediately
@@
This result confirms that the Gaussian curvature changes sign based on the value of $r$: for $r<\frac{3q^{2}}{2m}$, we find $02mr-3q^{2}<$ 0, making the Gaussian curvature $K>0$. Conversely, for $r>\frac{3q^{2}}{2m}$, we have $02mr-3q^{2}>$ 0, resulting in $K<0$.
To obtain the curvature for the $4$-dimensional case, we can derive the Riemann curvature tensor for the pseudo-Reissner-Nordström spacetime by applying a Wick rotation to the well-known Reissner-Nordström spacetime. This approach uses the established properties of the Reissner-Nordström solution, allowing us to modify the spacetime signature and adapt the curvature properties for the pseudo-Reissner-Nordström framework.
@@ 6.1. Pseudo-Reissner-Nordström spacetime with two horizons
The pseudo-Reissner-Nordström spacetime $M=S^{1}\times(0,\infty)$ with $m>q$ and two horizons at $r=r_{\pm}$ presents a particularly interesting configuration. These horizons divide $M$ into three distinct regions: $M_{3}=\{(\nu,r)\in M\mid 0<r<r_{-}\}$, $M_{2}=\{(\nu,r)\in M\mid r_{-}<r<r_{+}\}$, and $M_{1}=\{(\nu,r)\in M\mid r_{+}<r<\infty\}$. In analogy with the pseudo-Schwarzschild spacetime, the coordinate vector field $\partial_{r}$ is null, with $-\partial_{r}$ designated as future-pointing (indicating that $r$ decreases as one moves from past to future). The behavior of the vector field $\partial_{\nu}$ varies across regions: it is spacelike in $M_{1}$, timelike in $M_{2}$, and again spacelike in $M_{3}$.
@@
[…]
@@ 6.1.1. Non-chronal region $M_{2}$
Although the closed timelike curves (CTCs) in this case primarily arise from the periodicity of the $\nu$-coordinate, the tilting of light cones also contributes to the formation of CTCs.
@@
The inner product of two future-pointing causal vectors is non-positive. We already defined $-\partial_{r}$ to be future-pointing, thus the claim follows directly from $0\geq g(\gamma^{\prime},-\partial_{r})=-\gamma_{1}^{\prime}\Longleftrightarrow\gamma_{1}^{\prime}\geq 0$. ∎
@@
Assume $\gamma(0)\in M_{1}$, then $\underset{<0}{\underbrace{\frac{2}{(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)}}}<\frac{\gamma_{1}^{\prime}}{\gamma_{2}^{\prime}}$ follows from the aforegoing Lemma. As we approach the Cauchy horizon $H_{+}$ at $r_{+}$, we have $(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)\longrightarrow 0$, and thus $\frac{2}{(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)}\longrightarrow-\infty$. Consequently, $-\infty<\frac{\gamma_{1}^{\prime}}{\gamma_{2}^{\prime}}$ as we approach $H_{+}$. The curve would remain in $M_{1}$ and not enter region $M_{2}$ only if $\gamma_{2}^{\prime}=0$, meaning that it would be tangent to $\partial_{\nu}$ and thus lightlike. However, since $\gamma$ is timelike and $\gamma_{2}^{\prime}\neq 0$, it follows that the curve must enter the region $M_{2}$.
@@ 6.2. Extremal pseudo-Reissner-Nordström spacetime
The extremal $2$-dimensional pseudo-Reissner-Nordström spacetime $M$ with $m=q$ has only one horizon at $r=m$, which separates $M$ into two regions: $M_{3}=\{(\nu,r)\in M\mid 0<r<m\}$ and $M_{2}=\{(\nu,r)\in M\mid m<r\}$. As before, we time-orient $M$ by requiring $-\partial_{r}$ to be future pointing. The metric degenerates to
@@
where for $r<\frac{3m}{2}$, we have $02mr-3m^{2}<$ 0, so the Gaussian curvature is $\frac{2mr-3m^{2}}{r^{4}}>0$, and for $r>\frac{3m}{2}$, the curvature becomes negative.
@@
For a causal future-pointing curve $\gamma:[0,1]\rightarrow M_{2}:\,s\longmapsto\gamma(s)$ we have the requirement $\gamma_{1}\geq 0$, and $\gamma_{2}^{\prime}\leq\underset{\leq 0}{\underbrace{-\gamma_{1}^{\prime}\cdot\frac{(m-r)^{2}}{r^{2}}\cdot\frac{1}{2}}}$ $\Longrightarrow\gamma_{2}^{\prime}\leq 0$. For $\gamma_{2}^{\prime}=0$ we have $00\leq\underset{\leq 0}{\underbrace{-\gamma_{1}^{\prime}\cdot\frac{(m-r)^{2}}{r^{2}}\cdot\frac{1}{2}}}\Longleftrightarrow\gamma_{1}^{\prime}=0\,\vee\,\frac{(m-r)^{2}}{r^{2}}=$ 0. Note that $\gamma_{1}^{\prime}=0$ implies, on the one hand, that $\gamma^{\prime}=\gamma_{1}^{\prime}\partial_{\nu}+\gamma_{2}^{\prime}\partial_{r}=0$, which means $\gamma^{\prime}$ is not future pointing. On the other hand, the solution to $\frac{(m-r)^{2}}{r^{2}}=0$ gives $r=m$, which is the horizon and is not part of region $M_{2}$. Hence, we conclude that $\frac{dr}{ds}=\gamma_{2}^{\prime}<0$, as claimed. ∎

### v3

@@ 5.1.2. Relation in the 44–dimensional case
In earlier work [18] we conjectured that the 44–dimensional pseudo-Schwarzschild spacetime and Misner space are isocausal. In the present paper, we prove this result and extend it to include the pseudo-Reissner-Nordström spacetime. The proof, given in Proposition 7.1, establishes that all three models are pairwise isocausal on their universal covers and on suitable causally regular regions of the compactified spacetimes, and provides explicit criteria for when this equivalence descends globally.
@@ 6. Pseudo-Reissner-Nordström spacetime
By performing a Wick rotation $\theta\rightarrow i\theta$ on the Reissner-Nordström metric, changing the range of $\theta$ from $0,\pi]$ to $0,+\infty]$, flipping the signature, and identifying $t+P$ with $t$, we obtain the manifold
@@
Similar to the pseudo-Schwarzschild manifold, the 11–sphere $S^{1}$ (with $0\leq t<t+\alpha$) is topologically equivalent to an interval with its endpoints identified. The surface $H_{r}^{2}$ is 22–dimensional and corresponds to the upper sheet of a two-sheeted spacelike hyperboloid. The coordinate $\theta$ can take any positive value, while $\phi$ is periodic and ranges from $0\leqslant\phi<2\pi$. The real constants $m$ and $q$ represent mass and electric charge, respectively, and are assumed to be positive. This new metric can also be derived by performing a Wick rotation and a signature change on the Schwarzschild metric, followed by the substitution
@@
The Reissner-Nordström spacetime is not a vacuum solution of Einstein’s field equations, as it contains matter in the form of an electromagnetic field. It is instead an electrovacuum solution of the Einstein-Maxwell equations, described by an energy-momentum tensor $T_{\mu\nu}$ sourced by the electromagnetic field (from a vector potential). The metric, with signature (+++)(-+++), satisfies the Einstein equations in the form $G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=\kappa T_{\mu\nu}$, where $T_{\mu\nu}$ is the energy-momentum tensor of the electromagnetic field.
Our construction of the pseudo-Reissner-Nordström spacetime begins by performing a Wick rotation on the angular coordinate, $\theta\rightarrow i\theta$, followed by a sign reversal of the metric tensor, $g^{\prime}_{\mu\nu}=-g_{\mu\nu}$. This transformation results in a hyperbolic geometry in the angular sector. While the Einstein tensor $G_{\mu\nu}$ remains invariant under this sign reversal, the energy-momentum tensor, which is constructed from the metric, changes sign such that $T^{\prime}_{\mu\nu}=-T_{\mu\nu}$. Consequently, the field equations for the new metric become $G_{\mu\nu}=\kappa T^{\prime}_{\mu\nu}=-\kappa T_{\mu\nu}$.
This resulting spacetime is no longer a solution of the standard Einstein-Maxwell equations. The gravitational source now corresponds to an energy-momentum tensor that violates the weak energy condition, as indicated by $T_{\mu\nu}k^{\mu}k^{\nu}<0$ for some timelike vector $k^{\mu}$. This violation implies the presence of a source of exotic matter with negative energy density. This situation differs significantly from the pseudo-Schwarzschild metric, which remains a valid vacuum solution because its original stress-energy tensor was zero and remains so after the analogous transformation.
This distinction is critical. While the Misner and pseudo-Schwarzschild spacetimes are both vacuum solutions, the pseudo-Reissner-Nordström spacetime is a non-vacuum solution that requires exotic matter. This non-vacuum character is what makes it a particularly surprising and significant member of the family of causality-violating spacetimes we are investigating.
@@
We regard $\theta$ and $\phi$ being constant and first compute the tortoise coordinate rr* by setting
@@
$t(r)$ $\int\frac{1}{(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)}dr$
$(2m^{2}-q^{2})\cdot\tan^{-1}(\frac{r-m}{\sqrt{q^{2}-m^{2}}})-m\cdot\log(-2mr-q^{2}+r^{2})-r+c,$
@@
$ds^{2}$ $(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)\,(d(\nu-r*))^{2}+(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sinh^{2}(\theta)d\phi^{2})$
$(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)\,(\underset{(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{-1}}{\underbrace{\frac{\partial(\nu-r*)}{\partial r}}}dr+\underset{1}{\underbrace{\frac{\partial(\nu-r*)}{\partial\nu}}}d\nu)^{2}$
$(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sinh^{2}(\theta)d\phi^{2})$
$(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)\,[(\frac{1}{(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{2}})dr^{2}-2\cdot 1\cdot\frac{1}{(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)}drd\nu+1d\nu^{2}]$
$(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sinh^{2}(\theta)d\phi^{2})$
$(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{-1}dr^{2}+2drd\nu-(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)d\nu^{2}$
$(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sinh^{2}(\theta)d\phi^{2})$
$(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)d\nu^{2}+2drd\nu+r^{2}(d\theta^{2}+\sinh^{2}(\theta)d\phi^{2})$.
@@
The Riemann curvature tensor for the 22–dimensional pseudo-Reissner-Nordström spacetime can be obtained using the same approach as in Section 4 for the pseudo-Schwarzschild case. The scalar curvature then follows immediately
@@
This result confirms that the Gaussian curvature changes sign based on the value of $r$: for $r<\frac{3q^{2}}{2m}$, we find $2mr-3q^{2}<0$, making the Gaussian curvature $K>0$. Conversely, for $r>\frac{3q^{2}}{2m}$, we have $2mr-3q^{2}>0$, resulting in $K<0$.
To obtain the curvature for the 44–dimensional case, we can derive the Riemann curvature tensor for the pseudo-Reissner-Nordström spacetime by applying a Wick rotation to the well-known Reissner-Nordström spacetime. This approach uses the established properties of the Reissner-Nordström solution, allowing us to modify the spacetime signature and adapt the curvature properties for the pseudo-Reissner-Nordström framework.
@@ 6.1. Pseudo-Reissner-Nordström spacetime with two horizons
The pseudo-Reissner-Nordström spacetime $M=S^{1}\times(0,\infty)$ with $m>q$ and two horizons at $r=r_{\pm}$ presents a particularly interesting configuration. These horizons divide $M$ into three distinct regions: $M_{3}=\{(\nu,r)\in M\mid 0<r<r_{-}\}$, $M_{2}=\{(\nu,r)\in M\mid r_{-}<r<r_{+}\}$, and $M_{1}=\{(\nu,r)\in M\mid r_{+}<r<\infty\}$. In analogy with the pseudo-Schwarzschild spacetime, the coordinate vector field $\partial_{r}$ is null, with $\partial_{r}$ designated as future-pointing (indicating that $r$ decreases as one moves from past to future). The behavior of the vector field $\partial_{\nu}$ varies across regions: it is spacelike in $M_{1}$, timelike in $M_{2}$, and again spacelike in $M_{3}$.
@@
where $R=2mr-q^{2}-r^{2}$, $a=-q^{2}$, $b=2m$ and c=1c=-1. This then takes the form of the integral $\nu(r)=2\int\frac{r^{2}}{R}dr=2(-r-m\ln R+(2m^{2}-q^{2})\int\frac{1}{R}dr)$, where $\triangle=4ac-b^{2}=4(q^{2}-m^{2})<0$, and
$\int\frac{1}{R}dr=\frac{1}{\sqrt{-\triangle}}\ln\frac{\sqrt{-\triangle}-(b+2cr)}{(b+2cr)+\sqrt{-\triangle}}=\frac{1}{\sqrt{-4(q^{2}-m^{2})}}\ln\frac{\sqrt{-4(q^{2}-m^{2})}-(2m-2r)}{(2m-2r)+\sqrt{-4(q^{2}-m^{2})}}.$
@@ 6.1.1. Non-chronal region $M_{2}$
Although the closed timelike curves (CTCs) in this case primarily arise from the periodicity of the $\nu$–coordinate, the tilting of light cones also contributes to the formation of CTCs.
@@
The inner product of two future-pointing causal vectors is non-positive. We already defined $\partial_{r}$ to be future-pointing, thus the claim follows directly from $0\geq g(\gamma^{\prime},-\partial_{r})=-\gamma_{1}^{\prime}\Longleftrightarrow\gamma_{1}^{\prime}\geq 0$. ∎
@@
Assume $\gamma(0)\in M_{1}$, then $\underset{<0}{\underbrace{\frac{2}{(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)}}}<\frac{\gamma_{1}^{\prime}}{\gamma_{2}^{\prime}}$ follows from the aforegoing Lemma. As we approach the Cauchy horizon $H_{+}$ at $r_{+}$, we have $(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)\longrightarrow 0$, and thus $\frac{2}{(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)}\longrightarrow$-$\infty$. Consequently, $\infty<\frac{\gamma_{1}^{\prime}}{\gamma_{2}^{\prime}}$ as we approach $H_{+}$. The curve would remain in $M_{1}$ and not enter region $M_{2}$ only if $\gamma_{2}^{\prime}=0$, meaning that it would be tangent to $\partial_{\nu}$ and thus lightlike. However, since $\gamma$ is timelike and $\gamma_{2}^{\prime}\neq 0$, it follows that the curve must enter the region $M_{2}$.
@@ 6.2. Extremal pseudo-Reissner-Nordström spacetime
The extremal 22–dimensional pseudo-Reissner-Nordström spacetime $M$ with $m=q$ has only one horizon at $r=m$, which separates $M$ into two regions: $M_{3}=\{(\nu,r)\in M\mid 0<r<m\}$ and $M_{2}=\{(\nu,r)\in M\mid m<r\}$. As before, we time-orient $M$ by requiring $\partial_{r}$ to be future pointing. The metric degenerates to
@@
where for $r<\frac{3m}{2}$, we have $2mr-3m^{2}<0$, so the Gaussian curvature is $\frac{2mr-3m^{2}}{r^{4}}>0$, and for $r>\frac{3m}{2}$, the curvature becomes negative.
@@
For a causal future-pointing curve $\gamma:[0,1]\rightarrow M_{2}:\,s\longmapsto\gamma(s)$ we have the requirement $\gamma_{1}\geq 0$, and $\gamma_{2}^{\prime}\leq\underset{\leq 0}{\underbrace{-\gamma_{1}^{\prime}\cdot\frac{(m-r)^{2}}{r^{2}}\cdot\frac{1}{2}}}$ $\Longrightarrow\gamma_{2}^{\prime}\leq 0$. For $\gamma_{2}^{\prime}=0$ we have $0\leq\underset{\leq 0}{\underbrace{-\gamma_{1}^{\prime}\cdot\frac{(m-r)^{2}}{r^{2}}\cdot\frac{1}{2}}}\Longleftrightarrow\gamma_{1}^{\prime}=0\,\vee\,\frac{(m-r)^{2}}{r^{2}}=0$. Note that $\gamma_{1}^{\prime}=0$ implies, on the one hand, that $\gamma^{\prime}=\gamma_{1}^{\prime}\partial_{\nu}+\gamma_{2}^{\prime}\partial_{r}=0$, which means $\gamma^{\prime}$ is not future pointing. On the other hand, the solution to $\frac{(m-r)^{2}}{r^{2}}=0$ gives $r=m$, which is the horizon and is not part of region $M_{2}$. Hence, we conclude that $\frac{dr}{ds}=\gamma_{2}^{\prime}<0$, as claimed. ∎